### v2

@@ 3.1 Experimental Setup
CheXagent $0.3321$ $0.1817$ $0.4526$ $0.3429$ $0.3273$
MedGemma $0.3827$ $0.1624$ $0.5648$ $0.5723$ $0.4205$
GPT-4o $0.3219$ $0.0928$ $0.4122$ $0.2127$ $0.2599$
GPT-4o w/ ReAct $0.3613$ $0.1221$ $0.5034$ $0.4619$ $0.3622$
GPT-4o w/ Workflow $0.4058$ $0.1617$ $0.5498$ $0.5351$ $0.4130$
RadAgents wo/ V-RAG $0.4128$ $0.1925$ $0.5147$ $0.5841$ $0.4260$
RadAgents $0.4426$ $0.2314$ $0.5758$ $0.6032$ 0.4632
[…]
CheXagent $0.2916$ $0.1318$ $0.4129$ $0.1825$ $0.2547$
MedGemma $0.2413$ $0.1189$ $0.4728$ $0.2416$ $0.2686$
GPT-4o $0.2237$ $0.1324$ $0.4635$ $0.3138$ $0.2833$
GPT-4o w/ ReAct $0.3521$ $0.1329$ $0.4731$ $0.3325$ $0.3226$
GPT-4o w/ Workflow $0.4080$ $0.1556$ $0.5187$ $0.3841$ $0.3653$
RadAgents wo/ V-RAG $0.4412$ $0.1826$ $0.5238$ $0.3821$ $0.3824$
RadAgents $0.4727$ $0.1829$ $0.5644$ $0.4527$ 0.4182
3.2 Existence and Attributes
@@
The VQA questions cover seven common findings in CXR: atelectasis, cardiomegaly, consolidation, edema, lung opacity, pleural effusion, and pneumothorax, derived from the standard test split of MS–CXR. Each image receives the following prompt (details are stated in the Appendix D):
@@
Table 3.1 shows that adding V-RAG to the agent improves all metrics over the ablation: +0.0298 CheXbert, +0.0389 RadGraph, +0.0611 RaTE, +0.0191 GREEN, raising Avg. from 0.4260 to 0.4632 (+0.0372, +8.7%). Among baselines, MedGemma is strongest (0.4205 Avg.) yet remains 10.2% below RadAgents. GPT-4o benefits from ReAct (+0.1023 Avg.) but still trails the full agent by 0.1010.
3.3 Comparison and Progression
@@
Given current image <image>, and previous image <image>, decide if [finding] is improving, stable, or worsening.
Figure 3 shows that RadAgents+V-RAG achieves the best overall accuracy on MS-CXR-T, surpassing the ablated agent and all LLM baselines. The ordering mirrors E&A, underscoring the value of retrieval over similar studies for longitudinal reasoning.
3.4 Report Generation
@@
We construct a MIMIC–CXR subset aligned with MS–CXR identities so that findings queried in VQA are represented in the corresponding reports. Prompts request generation of the Findings section, and all agents are activated by default. The prompt combines the template from Section 3.2 with a curated list of clinically significant findings, following prior work (tu2024towards; peng2025scaling); details appear in Appendix D.
Table 3.1 shows that, RadAgents attains the best report quality (Avg. 0.4182). V-RAG contributes +0.0335 Avg. (+8.7%), with the largest gain on GREEN (+0.0706), plus lifts on CheXbert (+0.0315) and RaTE (+0.0406). GPT-4o+ReAct improves over GPT-4o (+0.0393) but remains 0.0956 below RadAgents, while MedGemma drops on this task (0.2686 Avg.).
Figure 3: VQA results regarding progression.
3.5 Effectiveness of RadAgents Design
We evaluate the multi-agent design against a monolithic baseline in which a single LLM follows the ABCDE scheme (GPT-4o w/ Workflow), thereby separating improvements attributable to multi-agent coordination from those due to a structured workflow. Consistent with Table 3.1 and Figure 3, (i) introducing a workflow confers substantial gains over ReAct, and (ii) the multi-agent architecture further improves performance. These effects stem from contextual isolation within the workflow, which limits long-context drift in extended reasoning chains.
In addition, we validate the context verifier and conflict‐resolution modules. On our 1,147‐case dataset, the context verifier was triggered in 37.67% of instances. We observed a tool‐conflict rate of 32.78%, of which 78.99% were resolved correctly.
4 Discussion
In this work, we propose the first radiologist-like agentic system, which demonstrates superior performance in CXR interpretation. Further improvements are required for broader application: (1) the context verifier can only detect errors but cannot modify outputs, especially for visual evidence such as bounding boxes, so stronger supporting models are needed; (2) current experiments are limited to frontal-view images, as most existing tools are incompatible with lateral views; (3) prompting strategies could be further optimized; (4) the framework could be extended to additional modalities, such as CT and MRI; and (5) our current implementation relies on large proprietary models and multiple tools, making it costly and difficult to reproduce; exploring smaller, open-source models is a valuable direction.
Acknowledgments
Appendix A Demonstrations
A.1 Full reasoning trajectory for tracheal deviation detection
Figure 5 shows the reasoning trajectory for the query “Is the trachea midline?” The sequence proceeds as follows: (a) the orchestrator agent analyzes the query, activates the airway agent, and initiates actions; (b) the returned tool context is visualized; (c) the synthesizer agent integrates the context but issues a warning due to low-confidence segmentation results, triggering the context verifier—specifically, GPT-4o is invoked to re-judge and answer the query; and (d) synthesis is completed, producing the final answer with an associated confidence score.
A.2 Failure case in CTR calculation
Most measurement reasoning relies on segmentation or grounding tools for support. For example, measuring heart width directly involves counting the number of pixels between boundary points. However, due to the high cost of annotation, ground-truth datasets for training robust segmentation or grounding models are limited. Moreover, most organ segmentation datasets are curated from normal images; when abnormalities obscure the organs, segmentation often fails, leading to downstream measurement errors. Figure 4(b) illustrates the segmentation masks for a normal CXR and a CXR with effusion. In the effusion case, both heart width and thoracic width are measured incorrectly due to inappropriate masks. These results further emphasize the need for a context verifier and resolver.
Figure 4: CTR calculation with segmentation masks on (a) normal and (b) effusion cases. Inaccurate masks in the abnormal case lead to incorrect heart and thoracic width measurements.
Figure 5: RadAgents’ reasoning trajectory for tracheal deviation detection.
Appendix B Clinical Evaluation Metrics
Evaluating the quality of generated radiology reports is non-trivial. Early works adopted general-domain natural language processing metrics such as ROUGE (lin2004rouge) and BLEU (papineni2002bleu). While these metrics are widely used for text evaluation, they treat differences in wording the same as clinically significant errors, failing to reflect medical accuracy. To address this limitation, clinically informed evaluation metrics, such as CheXbert (smit2020combining), RadGraph (jain1radgraph), GREEN (ostmeier2024green), and RaTEScore (zhao2024ratescore), have been proposed to better assess clinical correctness and utility. CheXbert is based on multi-label classification results for 5 or 13 diseases (along with one extra “normal” label). RadGraph considers literal entity agreement considering the positive or negative context of each entity. GREEN judges recall and precision errors by LLM prompting. RaTEScore is inspired by RadGraph but less sensitive to phrasing by an F1-like computation which allows semantic matching between entities based on a cosine similarity. The metrics are computed using their official and standardized implementations: RadGraph-F1111https://pypi.org/project/radgraph/0.1.2/, CheXbert-F1222https://pypi.org/project/f1chexbert/, RaTE Score [^3] https://pypi.org/project/RaTEScore/0.5.0/, and GREEN [^4] https://pypi.org/project/green-score/0.0.8/.
Appendix C Dataset
Operating an agentic system is costly because multimodal multi-step reasoning entails many LLM API calls; accordingly, we did not conduct very large scale experiments. We evaluate on MS–CXR and MS–CXR–T, whose annotations enable verification of intermediate outputs and support future reinforcement learning to incentivize tool use with open source models. For report generation, we condition on the current study and a single prior frontal image, rather than the full imaging history, to limit context length and processing cost. Dataset statistics are shown in Table 6.
Dataset # Cases # Images per case Has prior?
MIMIC-CXR (subset) 181 2 Yes
MS-CXR (test set) 181 1 No
MS-CXR-T 785 2 Yes
Table 2: Details of datasets used in RadAgents.
Appendix D Prompting
The input to RadAgents includes not only the image and query but also optional clinical context, such as patient demographics, indication, acquisition technique, comparison studies, and examination details (e.g., view/projection and region). We denote this aggregated context as Context in the templates for comprehensive CXR scanning and for report generation.
{infobox}
Instruction: You are a helpful radiology assistant. Describe what lines, tubes, and devices are present and each of their locations. Describe if pneumothorax is present; if present, describe size on each side. Describe if pleural eusion is present; if present, describe amount on each side. Describe if lung opacity (atelectasis, brosis, consolidation, inltrate, lung mass, pneumonia, pulmonary edema) is present; if present, describe kinds and locations. Describe the cardiac silhouee size. Describe the width and contours of the mediastinum. Describe if hilar enlargement is present; if enlarged, describe side. Describe what fractures or other skeletal abnormalities are present.
Given the [view] X-ray image(s) <images>,
Q: Describe the findings in the image following the Instructions, and Context: {context}.
Appendix E Reasoning modes and agentic workflows
Why interleaved reasoning?
Classical vision and language systems compress the image once and then reason only in text. This single pass approach is ill suited to radiology, where clinicians iteratively survey the study, escalate suspicious findings to measurement, revisit earlier impressions as new evidence appears, and compare with priors. RadAgents operationalizes this practice by interleaving perception and reasoning: at intermediate points the system inspects additional visual evidence (for example, cropped regions or segmentation overlays) and updates its hypothesis before proceeding. A simple protocol is a two phase visual chain of thought (VCoT): first, a pure visual assessment that answers the question and cites observable evidence without tool outputs; second, an evidence validation step that reveals tool conclusions (for example, a measurement or a mask) and records agreement, disagreement, or uncertainty with a recalibrated confidence. This improves transparency and reduces anchoring on imperfect tools. Figure 6 illustrates how interleaved quantitative evidence aids enlarged heart detection, whereas a simple grounding guided approach fails.
(M1) Measurement
Goal. Provide objective and reproducible judgments for geometry constrained findings.
When. Explicit measurement requests or size abnormalities suggested by a sweep.
Tools. Segmentation yields organ masks and derived metrics; the cardiothoracic ratio (CTR) uses maximal cardiac width and thoracic width from bilateral lung extents, with projection recorded (PA versus AP) and raw pixel widths logged.
Evidence. Numeric values with projection, overlay visuals of the masks used, and brief caveats.
(M2) Localization
Goal. Localize small or subtle targets that benefit from high resolution crops.
When. Explicit localization queries or equivocal global signals.
Tools. Grounding proposes bounding boxes; segmentation constrains search when organ context matters; crop policies adapt to target scale.
Evidence. Boxes with confidence, region thumbnails, and landmark distances when relevant.
(M3) Characterization
Goal. Describe texture, morphology, distribution, and severity of parenchymal and pleural findings.
When. Opacity related queries or when a sweep suggests edema, atelectasis, pneumonia, or fibrosis.
Tools. Zone prioritization by a classifier, lung masks to focus attention, region crops from suspicious zones, VQA to standardize descriptors, and VCoT to justify labels.
Evidence. Pattern and distribution labels with severity, plus representative crops tied to the cited features.
(M4) Relational and comparative reasoning
Goal. Determine improvement, worsening, or stability relative to a prior study, as well as the relationships between each finding like complication.
When. Explicit comparison or association requests or any mention of a prior.
Tools. Temporal alignment of regions, adjustment for projection or view differences when possible, signed change statements in VCoT, and measurement deltas for quantitative targets.
Evidence. Paired crops and a concise comparison statement.
(M5) Differential Diagnosis
Goal. Synthesize evidence across modes to issue a patient level conclusion and differential with calibrated confidence.
When. After upstream modes have supplied sufficient evidence or when the query explicitly seeks a diagnosis.
Tools. The synthesizer aggregates structured findings, checks consistency, reconciles conflicts, and maps evidence to diagnostic statements aligned with clinical guidance.
Evidence. A short justification that links key measurements, locations, and patterns to the final conclusion, plus uncertainty notes when appropriate.
Appendix F Details of V-RAG
@@
Table 3: Retrieval quality at different $k$ values, showing the trade-off between helpful and harmful rates.
$k$ Helpful-rate Harmful-rate
1 0.48 0.09
3 0.62 0.14
5 0.65 0.20
7 0.66 0.25
9 0.67 0.28
11 0.67 0.31
@@
We study how retrieval quality changes with the number of retrieved studies $k$ with the sampled 100 cases from the MS-CXR test set used in this study. For each setting, we compute two metrics: helpful-rate, the percentage of retrieved studies that improve the answer, and harmful-rate, the percentage that hurt the answer. As shown in Table 3, As k increases, the harmful rate grows more quickly than the helpful, e.g., the helpful-rate increases from 0.48 at $k=1$ to 0.65 at $k=5$, while the harmful-rate also rises from 0.09 to 0.20. We hypothesize this is due to longer contexts imposing a heavier reasoning burden and increasing hallucination, consistent with prior LLM findings. This illustrates a trade-off: retrieving more studies provides greater chances of including helpful evidence but also increases the risk of introducing misleading content. To balance these effects, we choose $k=3$ by default, achieving a helpful-rate of 0.62 with a moderate harmful-rate of 0.14.
@@
This is the $i$-th similar image and its report for your reference. [Reference]i ... According to the query image and the references, [Question] [Query Image]. where each reference is denoted as $(I_{i},R_{i})$.

### v3

@@ 4 Discussion
Appendix A Pathology-aware Skill Set (Examples)
  • Lines, Tubes, and Devices.
    Goal: Identify device type, localize position, and verify termination points (e.g., ET tube depth).
    Workflow: (1) Detection: Use BiomedParser with prompts (e.g., ”Endotracheal tube”, ”CVC”) to generate device masks. (2) Anatomical Localization: Overlay device masks with anatomical masks (Trachea, Carina, SVC) from CXAS. (3) Verification: Use Python scripts to measure distances (e.g., ET tip to Carina) or check containment (e.g., NG tube tip inside Stomach mask). Malposition is flagged based on geometric thresholds.
  • Pneumothorax.
    Goal: Confirm presence, laterality, and estimate size.
    Workflow: (1) Screening: Use DenseNet-121 to filter negative cases based on probability thresholds. (2) Localization: Deploy MAIRA-2 with prompts like ”Pneumothorax” or ”Pleural line” to generate bounding boxes. (3) Side Determination: Map bounding box centroids to Left/Right Lung masks from CXAS. (4) Characterization: Query MedGemma (VQA) with cropped regions to estimate size (small/moderate/large) and check for tension physiology.
  • Pleural Effusion.
    Goal: Detect presence, laterality, and quantify volume.
    Workflow: (1) Segmentation: Use CXAS to segment Lung fields and Costophrenic Angles. (2) Detection: Use BiomedParser to segment ”Fluid” or ”Effusion” regions. (3) Geometric Measurement: Calculate the vertical height of the fluid mask relative to the total lung height. (4) Quantification: Apply rule-based thresholds (e.g., <1/4<1/4 lung height = Small; >1/2>1/2 = Large) and check for costophrenic angle blunting.
  • Lung Opacity (Pneumonia, Mass, Atelectasis, etc.).
    Goal: Differentiate pathology type (texture) and precise anatomical location.
    Workflow: (1) Classification: Use DenseNet-121 to obtain probability distributions for various opacity types. (2) Localization: Use MAIRA-2 to generate bounding boxes for high-probability findings. (3) Anatomy Mapping: Overlay bounding boxes with CXAS lung lobe masks to assign location (e.g., RUL, LLL). (4) Texture Analysis: Use MedGemma (VQA) to describe pixel patterns (e.g., ”patchy/fluffy” for pneumonia vs. ”linear/plate-like” for atelectasis) to confirm diagnosis.
  • Cardiac Silhouette.
    Goal: Assess heart size and calculate Cardiothoracic Ratio (CTR).
    Workflow: (1) Segmentation: Use CXAS to segment the Heart and Thoracic Cavity/Lungs. (2) Calculation: Execute Python scripts to measure the maximum horizontal cardiac width and thoracic width. (3) Thresholding: Calculate CTR (Cardiac Width / Thoracic Width); if >0.5>0.5, classify as Cardiomegaly/Enlarged.
  • Mediastinum and Hilar Regions.
    Goal: Evaluate mediastinal width/contours and hilar enlargement.
    Workflow: (1) Segmentation: Use CXAS to segment the Mediastinum and Aortic Knob. (2) Measurement: Measure superior mediastinal width via Python script (threshold >8>8 cm for widening). (3) Hilar Assessment: Use MAIRA-2 to detect ”Hilar enlargement” or ”Mass”; if detected, use VQA to characterize potential lymphadenopathy or vascular prominence.
  • Skeletal Abnormalities.
    Goal: Detect and localize fractures or bone lesions.
    Workflow: (1) Detection: Use BiomedParser to parse ”Fracture” or ”Bone lesion” into masks. (2) Localization: Intersect finding masks with CXAS skeletal masks (Ribs, Clavicles, Spine) to identify specific bones (e.g., ”Right 6th Rib”). (3) Confirmation: Use VQA on the ROI to verify cortical disruption and rule out false positives from vascular overlap.
Appendix B Anatomy-aware Skill Set (Examples)
  • Cardiomegaly (Cardiothoracic Ratio, CTR). Given heart and bilateral lung masks, we first identify the axial row where the heart mask attains its maximal width. On this row, we extract the leftmost and rightmost pixels of the heart mask to obtain the cardiac width, and the leftmost and rightmost pixels of the union of the left- and right-lung masks to obtain the thoracic width. The cardiothoracic ratio is then
    $\mathrm{CTR}=\frac{\text{cardiac width}}{\text{thoracic width}}.$
    Comparing CTR to a view-specific clinical threshold (e.g., $\mathrm{CTR}>0.5$ on PA view) yields a binary cardiomegaly label.
  • Carina Angle. From the trachea bifurcation mask we compute a concave hull and locate the lowest point where the trachea splits, defining the carina point. At $10\%$, $20\%$, and $30\%$ of the mask height below the carina, we find the leftmost and rightmost side points and average them to estimate the centerlines of the left and right main bronchi. The carina angle is defined as the interior angle between the two vectors from the carina point to the left and right bronchus centerline points, respectively. This angle is compared against a normal reference range to determine whether the carina angle is abnormal.
  • Tracheal Deviation. Using trachea, trachea bifurcation, and vertebrae (T1–T7) masks, we first suppress the carina region within the trachea mask, retaining the upper trachea. From the vertebral masks, we extract posterior spinous-process points that lie within the vertical extent of the trachea mask. For each such vertebral level, we compare the horizontal position of the trachea to the corresponding spinous-process point and assign a local label (left, right, centered). The final tracheal deviation label is obtained via majority voting across levels: predominant left labels indicate left deviation, predominant right labels indicate right deviation, and balanced or centered labels indicate no deviation.
  • Mediastinal Widening. From upper mediastinum and bilateral lung masks, we find the row where the mediastinum mask is widest. On this row, we extract the leftmost and rightmost mediastinum pixels to obtain the mediastinum width and the leftmost and rightmost pixels of the combined lung masks to obtain the thoracic width at the same level. The mediastinum-to-thoracic width ratio
    $\text{Ratio}=\frac{\text{mediastinum width}}{\text{thoracic width}}$
    is compared against an expert-defined threshold to determine the presence of mediastinal widening.
  • Aortic Knob Enlargement. Using aortic arch, descending aorta, trachea, and trachea bifurcation masks, we first localize the aortic knob with the aortic arch mask. The trachea mask, restricted to the vertical range of the aortic arch and excluding the bifurcation, provides the starting point of the knob measurement: Point A is the average $x$-coordinate of the left tracheal wall in this range. From the upper $30\%$ of the descending aorta mask, we extract the innermost $x$-coordinate (Point B) and the outermost $x$-coordinate on the same row (Point C). Point D is the outermost $x$-coordinate of the aortic arch mask. The aortic knob width is defined as the horizontal distance from Point A to the farthest of Point C or Point D. The tracheal width is computed as the median horizontal width of the trachea mask (excluding the bifurcation). The diagnostic index is
    $\text{Ratio}=\frac{\text{aortic knob width}}{\text{tracheal width}},$
    and aortic knob enlargement is declared when this ratio exceeds an expert-defined threshold.
  • Ascending Aorta Enlargement. With ascending aorta, heart, trachea, and trachea bifurcation masks, we first identify the most right-sided point of the heart mask (Point A) and the most right-sided point of the trachea mask excluding the bifurcation (Point B), which approximates the inner boundary of the right lung. We construct a reference line connecting Points A and B and then isolate the portion of the ascending aorta mask that extends beyond this line toward the right. Let $A_{\text{beyond}}$ be the area of the ascending aorta beyond the reference line and $A_{\text{total}}$ the total ascending aorta area. The diagnostic index is
    $\text{Ratio}=\frac{A_{\text{beyond}}}{A_{\text{total}}},$
    and if this ratio exceeds an expert-defined threshold, the ascending aorta is labeled enlarged.
  • Descending Aorta Enlargement. From the descending aorta, trachea, and trachea bifurcation masks (with the heart mask used to define the thoracic segment), we retain only the descending aorta portion lying above the inferior border of the heart, corresponding to the thoracic aorta. Within this region, we identify the row where the aortic width is maximal and extract the leftmost and rightmost $x$-coordinates at this row, yielding the maximum thoracic descending aorta width. From the trachea mask, we compute the median tracheal width across rows and record the corresponding left/right $x$-coordinates at the median row. The diagnostic index is
    $\text{Ratio}=\frac{\text{descending aorta width}}{\text{tracheal width}},$
    and descending aorta enlargement is indicated when this ratio exceeds an expert-defined threshold.
Appendix C Vanilla-MedRAX-comparable experiments using GPT-4o
C.1 Experimental Setup
Here we use GPT-4o as the core LLM engine instead of Qwen3-VL, and, to make our system comparable to vanilla MedRAX, we adopt the same tool set. Considering the cost of API calls, we perform our evaluation on smaller-scale datasets that are less structurally complex than ChestAgentBench but exhibit increasing reasoning complexity: 181 MS-CXR (boecking2022making) VQA cases for existence and attribute (E&A) queries about abnormalities, 785 VQA cases for comparison and progression (C&P) of abnormalities, and 181 MIMIC-CXR two-view frontal report-generation cases.
C.2 Results
Existence and attributes.
@@ C.2 Results
The VQA questions cover seven common findings in CXR: atelectasis, cardiomegaly, consolidation, edema, lung opacity, pleural effusion, and pneumothorax, derived from the standard test split of MS–CXR. Each image receives the following prompt:
@@
Comparison and progression.
@@
Given current image <image>, and previous image <image>, decide if [finding] is
improving, stable, or worsening.
Report generation.
@@
We construct a MIMIC–CXR subset aligned with MS–CXR identities so that findings queried in VQA are represented in the corresponding reports. Prompts request generation of the Findings section, and all agents are activated by default.
Table 2 summarizes performance across the three evaluation settings. RadAgents with V-RAG achieves the best results on all tasks. Relative to the ablated agent without V-RAG, it improves GREEN on VQA (E&A) from 0.5841 to 0.6032 (+0.0191, $\sim$3.3%), raises VQA (C&P) accuracy from 48.0% to 50.9% (+2.9 points, $\sim$6.0%), and boosts report-generation GREEN from 0.3821 to 0.4527 (+0.0706, $\sim$18.5%). Augmenting GPT-4o with ReAct and our workflow also yields consistent gains over the base LLM—for example, GREEN on VQA (E&A) increases from 0.2127 (GPT-4o) to 0.4619 (GPT-4o+ReAct) and 0.5351 (GPT-4o+Workflow), while VQA (C&P) accuracy rises from 18.2% to 41.9% and 45.5%, respectively. Nonetheless, even the strongest GPT-4o+Workflow baseline remains substantially behind RadAgents, trailing by 0.0681 GREEN on VQA (E&A), 5.4 accuracy points on VQA (C&P), and 0.0686 GREEN on report generation. Overall, the ranking of methods is largely consistent across tasks, suggesting that the benefits of structured tool use and V-RAG transfer from explanation-style VQA to longitudinal comparison questions and free-text report generation.
Table 2: Performance comparison between different setting when using gpt-4o as the core LLM engine in RadAgents.
Method VQA (E&A) (GREEN) VQA (C&P) (Acc. %) Report Generation (GREEN)
CheXagent 34.3 34.1 18.3
GPT-4o 21.3 18.2 31.4
GPT-4o w/ ReAct 46.2 41.9 33.3
GPT-4o w/ Workflow 53.5 45.5 38.4
RadAgents w/o V-RAG 58.4 48.0 38.2
RadAgents 60.3 50.9 45.3
Appendix D Details of V-RAG
@@
We study how retrieval quality changes with the number of retrieved studies $k$ with the sampled 100 cases from the MS-CXR test set used in this study. For each setting, we compute two metrics: helpful-rate, the percentage of retrieved studies that improve the answer, and harmful-rate, the percentage that hurt the answer. As shown in Figure 7, As k increases, the harmful rate grows more quickly than the helpful, e.g., the helpful-rate increases from 0.48 at $k=1$ to 0.65 at $k=5$, while the harmful-rate also rises from 0.09 to 0.20. We hypothesize this is due to longer contexts imposing a heavier reasoning burden and increasing hallucination, consistent with prior LLM findings. This illustrates a trade-off: retrieving more studies provides greater chances of including helpful evidence but also increases the risk of introducing misleading content. To balance these effects, we choose $k=3$ by default, achieving a helpful-rate of 0.62 with a moderate harmful-rate of 0.14.
@@
This is the $i$-th similar image and its report for your reference. [Reference]i
... According to the query image and the references, [Question] [Query Image].
where each reference is denoted as $(I_{i},R_{i})$.